Case B US UK Wheat 4 1 Cloth 3 2 Suppose that in Case B in...


Case B

US UK
Wheat 4 1
Cloth 3 2

Suppose that in Case B in the above tables, the US exchanges 4W for 4C with
the UK.
a. How much does the US gain in terms of C?
b. How much does the UK gain in terms of C?
c. What is the range for mutually beneficial trade?
d. How much would each nation gain in terms of C if they exchanged 4W for
6C instead?
Answer #1

Answer

(a)

We can see from above table that with given input US can either produce 4 Wheat(i.e. 4W) or 3 Cloth(i.e. 3C).

Hence, In order to produce 3C US have to sacrifice 4W. Here In exchange of 4W he gets 4C.

Hence US will gain 4C - 3C = 1C

Hence US will gain 1C

(b)

We can see from above table that with given input UK can either produce 1 Wheat(i.e. 1W) or 2 Cloth(i.e. 2C).

Hence, In order to produce 4C UK have to sacrifice 1*2W = 2W. Here In exchange of 4C he gets 4W.

hence UK will gain 4W - 2W = 2W and as disusses above 2W is worth 2*2C = 4C

Hence UK will gain 4C

(c)

As they both are gaining when US exports wheat and UK exports Cloth hence US will specialize in Wheat and UK will specialize in cloth.

We can see from above table that with given input US can either produce 4 Wheat(i.e. 4W) or 3 Cloth(i.e. 3C). Hence US will gain if for 1 wheat he gets more than (3/4)C and We can see from above table that with given input UK can either produce 1 Wheat(i.e. 1W) or 2 Cloth(i.e. 2C). Hence US will gain if for 1C he gets more than (1/2)W

hence they both will gain if

1W > (3/4)C and 1C > (1/2)W =>2C > 1W

=> (3/4)C < 1W < 2C --------------------Mutually beneficial trade

(d)

We can see from above table that with given input US can either produce 4 Wheat(i.e. 4W) or 3 Cloth(i.e. 3C).

Hence, In order to produce 3C US have to sacrifice 4W. Here In exchange of 4W he gets 6C.

Hence US will gain 6C - 3C = 3C

Hence US will gain 3C

We can see from above table that with given input UK can either produce 1 Wheat(i.e. 1W) or 2 Cloth(i.e. 2C).

Hence, In order to produce 6C UK have to sacrifice 1*3W = 3W. Here In exchange of 6C he gets 4W.  

hence UK will gain 4W - 3W = 1W and as discussed above 1W is worth 1*2C = 2C

Hence UK will gain 2C